# TAMS: Translation-Assisted Morphological Segmentation

### v1

@@ 5.3 Model and Hyperparameters
We treat translation incorporation strategy as a hyperparameter and test all combinations on the Tsez development set to determine the optimal approach.
@@
We take the average whole-word accuracy for configuration across all training split sizes on the Tsez development set using the CLS-None strategy. The full results of this search are shown in Table 1. From this search, we find that the overall highest performing strategy configuration is Init-State in the encoder and Concat-Half in the decoder. This configuration outperforms the baseline by an average of 1.58% and is the top performer on the 100 and 500 train sample settings. Only in the highest train data setting does the baseline outperform our top configuration.
@@
For the 100 train sample setting, Concat in the encoder and Concat-Half in the decoder, the second highest performing configuration overall, was the top configuration. With this in mind, we perform a second search over the CLS strategies on these two configurations, shown in Table 2. From this, we find that the overall highest performing configuration is Init-State in the encoder, Concat-Half in the decoder and CLS-Concat as our CLS Strategy; we call this model TAMS.
@@
6 Results with Final Model Configuration
Tsez Lezgi Arapaho
Train Limit Metrics TAMS Baseline TAMS Baseline TAMS Baseline
Accuracy 23.24 22.65 27.91 24.03 12.34 9.82
n = 100 F1 43.31 41.24 41.92 38.32 28.61 23.84
ED 3861 4065 808 902 37024 37719
Accuracy 35.39 31.05 33.25 34.71 19.65 18.08
n = 250 F1 57.14 54.74 51.06 53.00 41.36 40.61
ED 3156 3442 691 715 29022 30344
Accuracy 44.52 43.11 40.78 41.99 30.59 31.53
n = 500 F1 64.43 63.43 55.66 57.42 51.91 53.02
ED 2427 2655 636 639 21339 21261
Accuracy 80.78 82.60 46.84 44.66 67.72 67.08
n = all F1 89.52 90.44 62.48 60.75 81.62 81.11
ED 701 652 532 568 9899 10495
Table 3: Final results per language: Performance on all languages’ test sets on silver-aligned data.
Train Limit TAMS (abs/$\Delta$) Baseline
n = 100 21.16 (+2.33) 18.83
n = 250 29.43 (+1.48) 27.95
n = 500 38.63 (-0.25) 38.88
Full Train Split 65.11 (+0.33) 64.78
Table 4: Final results, all languages: Average accuracy (%) by training split size across all languages.
Model Metrics Gold Awesome-Gold
Accuracy 53.60 54.60
TAMS F1 67.98 67.89
ED 450 443
Baseline Accuracy 55.40
F1 70.19
ED 435
Table 5: Manual vs. automatic alignment: Performance on Tsez’s gold-aligned test split (n=500).
As described above, we perform search over the Tsez development set to determine the overall highest performing model, which we call TAMS. This configuration consists of Init-State in the encoder, Concat-Half in the decoder, and CLS-Concat as the CLS strategy.
6.1 Test Languages
We apply our final model configuration to the automatically aligned test sets in each of our three languages: Tsez (development language), Lezgi (in-family test language), and Arapaho (out-of-family test language.
Results are shown in Table 3. In most cases, we see that edit distance, F1, and accuracy are in agreement, so we focus on accuracy as our main evaluation metric. We report the average accuracy across languages for each training data setting in Table 4. We find that on average our model outperforms the baseline in every train data setting except for n=500. In general, performance gains are highest on the lower-resource settings, while on the higher-resource settings, performance improvements are slight if present at all. In the n=100 setting, TAMS outperforms the baseline by an average of 2.33%, suggesting that our model is most beneficial in truly low-data settings. Even for Lezgi, the language for which TAMS has the most varied results, on the 100 sample train-split TAMS yields a 3.88% increase in accuracy over the baseline. We also see consistent gains on Arapaho in all data settings except for n=500 which indicates that with further work, our model could be useful for polysynthetic languages. This is particularly exciting considering the relative difficulty of segmenting polysynthetic languages. However, TAMS’ performance on Lezgi is very uneven despite Lezgi’s linguistic similarity to Tsez, which indicates that the exact translation incorporation configuration we use in TAMS may not be fully extensible to other languages even if they are morphologically similar, and one may need to tune the translation strategy to fit their particular data setting.
6.2 Manual vs. Automatic Alignment
We additionally train a TAMS model on gold-aligned Tsez data (Gold) and compare it to a a TAMS model trained on equivalent data aligned with awesome-align (Awesome-Gold). We evaluate on gold-aligned data (note: this data split is distinct from the one in §6.1). We report the performance metrics of each model in Table 5. We find that the TAMS model trained with Awesome-Gold data actually outperforms its Gold counterpart on every metric except for F1. This suggests that gold alignment data might not be necessary to see the best possible performance from TAMS. This is a positive result because automatic alignment is much less costly than expert alignment and makes training a TAMS model more feasible for a broader range of languages. That said, the baseline outperforms both TAMS models, which complicates our analysis. Another potential issue is awesome-align’s use of mBERT, whose performance varies across languages Wu and Dredze (2020). Further work experimenting with different more languages and data resource levels is necessary to be more conclusive about gold vs. automatic alignment.
@@ 7 Conclusion
We present a novel method for incorporating information from translations into a morphological segmentation model to support low-resource canonical segmentation. Using Tsez as a development language, we determine our best-performing model (TAMS), which uses a fixed-length representation of the translation in two ways: to initialize the hidden state in the encoder (Init-State) and to concatenate to the input at each time step in the decoder (Concat-Half). Our model is most beneficial in the super low-resource setting (n=100), where it outperforms the baseline by 2.33% on average across three morphologically complex languages. And although we only tune our model on the Tsez development set, we also see impressive performance gains in all but one training split for Arapaho, a typologically and morphologically distinct polysynthetic language. This promising result suggests that TAMS could be beneficial for a wide range of languages. However, our results are more mixed in higher-resource settings which indicates that there is still more work to be done to determine whether translations are a valuable addition to canonical segmentation models in cases where more data is available. Overall, canonical segmentation for morphologically complex languages remains a challenging task, but we believe that this work indicates that translations should be explored further as an additional data resource.
There are several avenues for future work we wish to highlight. A first possible improvement strategy could be to experiment with providing more explicit information instead of or in addition to translations, such as the POS tags of the aligned English words. Second, it would be interesting to see whether our results can be reproduced on other languages and with other PLMs. Third, there may be other ways to use translation-based representations with LSTMs.
@@
Due to data availability, we experimented only on two language families, Northeast Caucasian and Algonquian, but ideally we would have tested on more language families. We cannot concretely say that our models would perform equivalently on a more diverse set of languages. Another limitation was in the exhaustiveness of our hyperparameter search. Ideally, we would have searched each possible CLS-token strategy with each possible configuration of translation incorporation strategies but we were unable to due to the GPU hours that would have been required.
@@ 8 Acknowledgements
Thanks to Bernard Comrie for helping us out with the Tsez dataset.
@@ A.1 High-level Directives
  • Align English prepositions with the Tsez word with the equivalent case marker
  • If some grammatical information is expressed in one language but not the other, behave as if it were expressed in the corresponding phrase. In such cases, the extra word (the, for example) should be aligned to the head of the corresponding phrase.
@@ A.2 Lower-level Directives
  • Pronominal subjects that are not expressed in Tsez and are expressed in English should have the English subject aligned to the head predicate in Tsez
@@
  • If a quotative verb like “say” is used a variable amount of times in one language compared to the other, then align all instances of the quotative verbs together, so long as the quoted material they’re all referring to is identical.
  • Always align the expletive there with the existential verb in Tsez. And if there is an adverbial there-equivalent in Tsez, do not align it with anything in English unless there really is an adverbial, non-expletive there or similar in English
@@ Appendix B Hyperparameter Tuning
We conduct hyperparameter tuning for the baseline Tsez models without translation using random search with the full training set of 6572. We then took the top two architectures and performed a hyper-parameter sweep with 100 and 500 training samples to simulate a lower-resource setting. The top performing models for each architecture and training set size are outlined in Tables 8, 7, 6. The search space of architecture specific hyperparameters is outlined in Table 9 and Table 10 and the search space of optimization parameters is outlined in Table 11. All models used Adam optimization. We report whole-word accuracy on the development set.

### v2

@@ 6.3 Supplementary Analysis of Translation Incorporation Strategies
In this section we provide details of the search that led to the final model configuration investigated above. We treat translation incorporation strategy as a hyperparameter and test all combinations on the Tsez development set to determine the optimal approach.
@@
We take the average whole-word accuracy for configuration across all training split sizes on the Tsez development set using the CLS-None strategy. The full results of this search are shown in Table 4. From this search, we find that the overall highest performing strategy configuration is Init-State in the encoder and Concat-Half in the decoder. This configuration outperforms the baseline by an average of 1.58 percentage points and is the top performer on the 100 and 500 train sample settings. Only in the highest train data setting does the baseline outperform our top configuration.
@@
For the 100 train sample setting, Concat in the encoder and Concat-Half in the decoder, the second highest performing configuration overall, was the top configuration. With this in mind, we perform a second search over the CLS strategies on these two configurations, shown in Table 5. From this, we find our final TAMS configuration: Init-State in the encoder, Concat-Half in the decoder and CLS-Concat as our CLS Strategy.
Tsez
@@ 7 Conclusion
We present a novel method for incorporating information from translations into a morphological segmentation model to support low-resource canonical segmentation. Using Tsez as a development language, we determine our best-performing model (TAMS), which uses a fixed-length representation of the translation in two ways: to initialize the hidden state in the encoder (Init-State) and to concatenate to the input at each time step in the decoder (Concat-Half). Our model is most beneficial in the super low-resource setting (n=100), where it outperforms the baseline by 1.99 percentage points on average across three morphologically complex languages. And although we only tune our model on the Tsez development set, we also see performance gains for Arapaho, a typologically and morphologically distinct polysynthetic language. This promising result suggests that TAMS could be beneficial for a wide range of languages. We believe TAMS will be the most beneficial in the language documentation setting, where extreme resource constraints are realistic and often expected.
The findings of our TAMS-CLS experiment are especially promising because they indicate that we may be able to see benefits from incorporating translations in canonical segmentation even without word-level alignments. This opens up TAMS to many more languages which may not have high-quality word-level alignments available.
However, our results are more mixed in higher-resource settings, which indicates that there is still more work to be done to determine whether translations are a valuable addition to canonical segmentation models in cases where more data is available. Overall, canonical segmentation for morphologically complex languages remains a challenging task, but we believe that this work indicates that translations should be explored further as an additional data resource. There are several avenues for future work we wish to highlight. A first possible improvement strategy could be to experiment with providing more explicit information instead of or in addition to translations, such as the POS tags of the aligned English words. Second, it would be interesting to see whether our results can be reproduced on other languages and with other PLMs. Third, there may be other ways to use translation-based representations with LSTMs.
@@
Due to data availability, we experimented only on two language families, Northeast Caucasian and Algonquian, but ideally we would have tested on more language families. We cannot concretely say that our models would perform equivalently on a more diverse set of languages. Another limitation was in the exhaustiveness of our hyperparameter search. Ideally, we would have searched each possible CLS token strategy with each possible configuration of translation incorporation strategies but we were unable to due to the GPU hours that would have been required.
@@ 8 Acknowledgements
We thank the anonymous reviewers for their useful suggestions and feedback, as well as the LECS and NALA Labs at the University of Colorado Boulder. We would like to extend our gratitude to Bernard Comrie and Sarah Moeller for their invaluable assistance with interpreting our datasets. Parts of this work were supported by the National Science Foundation under Grant No. 2149404, “CAREER: From One Language to Another.” Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. This work utilized the Blanca condo computing resource at the University of Colorado Boulder. Blanca is jointly funded by computing users and the University of Colorado Boulder.
@@ A.1 High-level Directives
  • Align English prepositions with the Tsez word with the equivalent case marker.
  • If some grammatical information is expressed in one language but not the other, behave as if it were expressed in the corresponding phrase. In such cases, the extra word (‘the’, for example) should be aligned to the head of the corresponding phrase.
@@ A.2 Lower-level Directives
  • Pronominal subjects that are not expressed in Tsez and are expressed in English should have the English subject aligned to the head predicate in Tsez.
@@
  • If a quotative verb like ‘say’ is used a variable amount of times in one language compared to the other, then align all instances of the quotative verbs together, so long as the quoted material they’re all referring to is identical.
  • Always align the expletive ‘there’ with the existential verb in Tsez. And if there is an adverbial ‘there’-equivalent in Tsez, do not align it with anything in English unless there really is an adverbial, non-expletive ‘there’ or similar in English
@@ Appendix B Hyperparameter Tuning
We conduct hyperparameter tuning for the baseline Tsez models without translation using random search with the full training set of 6572 words. We then took the top two architectures and performed a hyper-parameter sweep with 100 and 500 training samples to simulate a lower-resource setting. The top performing models for each architecture and training set size are outlined in Tables 8, 7, 6. The search space of architecture specific hyperparameters is outlined in Table 9 and Table 10 and the search space of optimization parameters is outlined in Table 11. All models used Adam optimization. We report whole-word accuracy on the development set.
@@
Appendix C Standard Deviation of TAMS Performance
Tsez Lezgi Arapaho Average
Train Metrics TAMS TAMS- Base TAMS TAMS- Base TAMS TAMS- Base TAMS TAMS- Base
Limit CLS CLS CLS CLS
Acc. 2.52 2.92 2.18 2.27 0.99 1.30 2.21 2.87 2.05 2.33 2.26 1.84
n = F1 2.87 2.83 2.23 3.57 3.70 4.10 3.04 3.73 3.43 3.16 3.42 3.25
100 ED 155 221 101 33 20 37 3006 3334 3727 1065 1192 1288
Acc. 1.62 1.77 0.93 1.96 2.07 1.51 1.30 1.75 0.96 1.63 1.86 1.11
n = F1 1.53 1.62 0.94 1.78 1.54 2.85 1.40 1.86 0.90 1.57 2.00 1.57
250 ED 117 152 165 23 31 29 1961 1292 1536 700 492 577
Acc. 1.54 1.78 0.89 2.37 1.92 2.15 1.59 1.69 1.28 1.83 1.80 1.44
n = F1 1.07 1.09 0.67 1.85 1.48 1.57 1.32 1.08 0.67 1.41 1.21 0.97
500 ED 110 120 64 27 24 31 688 967 918 275 370 338
Table 12: Standard deviation results per language: Standard deviation of performance metrics on all languages’ test sets (over 5 randomized test sets) on silver-aligned data. Metrics: Accuracy (Acc.), F1 score (F1), and Edit Distance (ED).
Tsez
Model Metrics Gold Awesome-Gold
Accuracy 1.40 2.48
TAMS F1 2.16 2.14
ED 33 42
Accuracy 1.48
TAMS-CLS F1 2.14
ED 33
Accuracy 1.98
Baseline F1 2.56
ED 48
Table 13: Standard deviation results on Tsez’s gold-aligned test split (n=100).